OpenStudy (anonymous):

5x-4=x^2+2 write each equation in standord form and determine (A), (B), and (C)

OpenStudy (anonymous):

x^2 - 5x + 6 = 0 a = 1, b= -5 and c = 6

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

w/come

OpenStudy (anonymous):

5X- x^2=6

OpenStudy (anonymous):

a quadratic standard form is defined as \[ax^2+bx+c=0\]. So transform the expression \[5x-4=x^2+2\] and you'll get: \[x^2-5x+2+4=0\]. Simplify this last equation and you should be a winner.

OpenStudy (anonymous):

@oojih u shud combine 2 and 4 to make 6

OpenStudy (anonymous):

@Harkirat, yeah, I don't want to solve the problem. I intentionally leave the last step for Chloebella to solve so that she understands how to solve this type of question.